Assuming you have $1500, can you use triangular arbitrage to generate a profit? Explain the strategy and order of transactions that are necessary for earning the profit. On the other hand, explain ‘why’ if you think profit is unattainable. Use the following information for making possible arbitrage strategy:-

You can buy a euro for 11 pesos. The bank will pay you 10 pesos for a euro. 

You can buy a U.S. dollar for .9 euros. The bank will pay you .8 Euros for a U.S. dollar. 

You can buy a U.S. dollar for 10 pesos.  The bank will pay you 9 pesos for a U.S. dollar.
Answer #1

If we have $1500 at present, then we would sell dollars and buy Euros.

Therefore, we will get = 1500*0.8 = 1200 Euros

Then we will buy pesos and sell 1200 Euros,

Hence we will get = 1200*10 = 12000 Pesos

We now have 12000 pesos which we will sell and buy dollars,

Therefore we will get = 12000/10 = $1200

Hence we can see that loss =1500-1200 = $300

Therefore, Aribtraging profit is unattainable.